OpenStudy (mony01):

Eliminate the parameter to find a Cartesian equation of the curve. x=t-1, y=t^(3)+1 -2<=t<=2

OpenStudy (anonymous):

t = x+1 so, y = (x+1)^3 + 1

OpenStudy (mony01):

oh thanks I thought we had to find the equation from the y=t^3+1 and get t^3=y-1

OpenStudy (mony01):

do i have to find new coordinates instead of using -2 and 2 do we find new numbers?

OpenStudy (dumbcow):

yes you should redefine the domain -2 <t< 2 becomes -3< x< 1 since x = t-1

OpenStudy (mony01):

thank you!!!

OpenStudy (dumbcow):

yw